# Analysis of anomalous contribution to galactic rotation curves due to stochastic spacetime

### v3

@@ Abstract
We analyze the model put forth by Ref. Oppenheim:2024rcp , in which it is claimed that “post-quantum classical gravity” (PQCG), a stochastic theory of gravity, leads to modified Newtonian dynamics (MOND) behavior on galactic scales that reproduces galactic rotation curves. We show that this analysis has three basic problems: (i) the equations of PQCG do not lead to a new large scale force of the form claimed in the paper, (ii) the form claimed is not of the MONDian form anyhow and so does not correspond to observed galactic dynamics, and (iii) the spectrum of fluctuations is very different from observations.
@@ I Introduction
An outstanding problem in modern physics is the successful unification of quantum mechanics and gravity. An interesting approach to this problem has been put forward in Refs. Oppenheim:2018igd ; Oppenheim:2023izn ; Layton:2023oud in which matter is treated quantum-mechanically, while gravity is treated classically; this is dubbed “post-quantum classical gravity” (PQCG). The coupling is such that gravity becomes effectively stochastic. Whether this framework truly leads to an internally consistent theory is beyond the scope of this note.
What is relevant to this note is the possibility of such theories leading to large scale testable predictions. Very interestingly, in Ref. Oppenheim:2024rcp it was claimed that PQCG indeed does so, namely that it leads to a new long range force between matter of the modified Newtonian dynamics (MOND) form; the form that can reproduce galactic rotation curves Milgrom:1983ca ; Milgrom:1983pn ; Milgrom:1983zz ; Bekenstein:1984tv . While it would be very interesting if the unification of gravity and quantum mechanics leads to such large scale effects. Here we point out that the PQCG theory of the form presented in Oppenheim:2024rcp does not in fact lead to anything like MONDian dynamics. Furthermore, we show that the theory has a very different spectrum of fluctuations than that observed. Whether some other variation of this framework improves upon this is beyond the scope of this note.
@@ II The Newtonian Limit of PQCG
The full PQCG theoretical framework is an interesting theory in which quantum dynamics of matter and classical dynamics of gravity are coupled together in a novel way. Nevertheless, the dynamics can be encoded in an action Oppenheim:2023izn . The full relativistic theory is somewhat complicated. However, for the purpose of studying galactic dynamics, we only need to pay attention to the low velocity limit of the theory as the characteristic speeds of gas, stars, and satellites in a galaxy are orders of magnitude slower than the speed of light (we set $c=1$). In Ref. Layton:2023oud , the low velocity limit of the theory is explained to be given by the following effective action
@@
If no specific boundary conditions are specified, then the most probable path simply minimizes the factor in brackets in Eq. (1). This gives the standard Newtonian potential $\Phi_{\mbox{\tiny{MPP}}}=\Phi_{N}$ obeying the Poisson equation
@@ III Solution of Modified Poisson Equation with Spherical Symmetry
For this section and the next, we work under the assumption that the fluctuations are small, leaving us with the MNE as the relevant equation for $\Phi$ (then we include fluctuations in Section V). The MNE is so similar to the standard Poisson equation for Newtonian gravity that we should expect very similar behavior with only subtle differences. However, Ref. Oppenheim:2024rcp claimed that there are dramatic differences. To unpack this, let us consider a localized source and consider the region outside of the source where $\rho=0$. In this vacuum region of space, the equation reduces to
@@
With the assumption of spherical symmetry, the general solution of this equation is
@@
where $\kappa_{-1,0,1,2}$ are constants in space, although it is not obvious they should be static in time (see below for more discussion). By comparing to the Newtonian theory, we can easily identify
@@ III.3 Another Derivation
Then the first term here was simply ignored apparently because it is merely a constant, leaving only the remaining $2GM\kappa_{1}/r^{2}$ term. By then taking the square root to recover the acceleration, it is stated that this gives the desired $1/r$ force law of MONDian dynamics. And the constant pre-factor $\kappa_{1}$ is to play the role of $a_{0}/2$ of MOND, where $a_{0}\sim 10^{-10}\,\mbox{m}/\mbox{s}^{2}$ is the critical acceleration in which Newton’s law transitions from $1/r^{2}$ to $1/r$.
However, this procedure is incorrect and is the second problem in the analysis. One cannot take a sum of two terms, $1/r^{2}$ and constant, square the sum, ignore the squared terms, only keep the cross term, then take the square root of the cross term to obtain the geometric mean $1/r$.
Instead the acceleration is Newton’s $1/r^{2}$ plus a constant. And there is no evidence that a constant correction helps to reproduce the observed galaxy rotation curves. Test particles do not respond to other objects in a way independent of their distance or mass.
@@ V.2 Full Distribution
Let us now examine in some detail the actual distribution. Often probability distributions can have interesting temporal dependence through the $dt$ integral, giving a rule for how to update the distribution from an initial time $t_{i}$ to a final time $t_{f}$.
@@
In fact as written the form presented is not well defined, as there are no time derivatives in this action. This means that in the path integral each moment in time is decoupled from the others. By breaking up the integral over time into a Riemann sum $\int_{t_{i}}^{t_{f}}\to\epsilon\sum_{i}^{f}$, where $\epsilon=dt$ is the time step, we see that the probability distribution factorizes and so all earlier times become unimportant. In the continuum $\epsilon=dt\to 0$ limit, we then have an un-normalized distribution, unless one sends the factor $\alpha\to\infty$ to compensate. If one does this, or if one simply introduces a hard cut off in time ($\epsilon$ remains finite), then one can normalize the distribution, but one should expect the distribution to jump around in time in an uncorrelated fashion. It could be that when one includes the contribution to the path integral from the (quantum) matter degrees of freedom, the situation is improved; we do not develop this issue further here. In any case, we shall proceed as is done in Ref. Oppenheim:2024rcp by ignoring the $dt$ integral and the temporal dependence. If there are no prior fixed boundary conditions, then the probability distribution for a fluctuation $\phi$ at any moment in time is the Gaussian distribution
@@
where $\alpha_{T}=\alpha\,\epsilon$ has units of length (we set $c=1$ here; if we reinstate factors of $c$, it has units of time4/length3).
@@ V.3 Power Spectrum
We note that since this is all derived in the Newtonian approximation, we only expect it to apply on sub-horizon scales, i.e., $k\gtrsim H_{0}$ the Hubble constant.
@@
$\langle\phi({\bf x})\,\phi({\bf y})\rangle$ $=$ $\int{d^{3}k\over(2\pi)^{3}}e^{-i{\bf k}\cdot({\bf x}-{\bf y})}P_{\phi}(k)$ (38)
@@
$\sigma_{k}\sim\sqrt{k^{3}\,P_{\phi}(k)\over 2\pi^{2}}={1\over 2\pi\sqrt{\alpha_{T}\,k}}$ (40)
@@ V.3.2 Large Scale Structure
Moreover, on sub-horizon scales, we can test if this spectrum of fluctuations is compatible with observations. The concordance model (CM) in cosmology with baryons, dark matter, and dark energy has a spectrum of fluctuations for $\Phi=\Phi_{N}$ (deterministic, but arising from inhomogeneous matter $\delta\rho_{m}=\bar{\rho}_{m}\,\delta_{m}$ drawn from some distribution through $\nabla^{2}\Phi_{N}/a_{s}^{2}=4\pi G\bar{\rho}_{m}\delta_{m}=3H^{2}\Omega_{m}^{2}\delta_{m}/2$) that is nicely compatible with observations. In linear theory, it takes on the form (for a review, see Ref. Hertzberg:2012qn )
@@
i.e., a flat spectrum. (In the CM, $P_{m,\mbox{\tiny{CM}}}$ is physically interpreted as the “matter power spectrum”, while in PQCG it does not directly have this interpretation as the fluctuations $\phi$ have no source. But what we observe are the effects of the gravitational potential, and so this factor of ($4k^{4}/(9\Omega_{m,0}^{2}H_{0}^{4})$ can be viewed as just a convenient re-scaling.)
@@
We see that the prediction from PQCQ of a flat power spectrum in the $P_{m}(k)$ variable is very different from the observed power in Figure 1. By comparing Eqs. (37) and (46), if $\alpha_{T}\sim 0.01/H_{0}$ as chosen in Ref. Oppenheim:2024rcp , then the power in PQCG is $\approx 10$ orders of magnitude too large on the scale of the horizon. And furthermore, it remains far too large even for modes that are several orders of magnitude within the horizon. This spectrum is clearly ruled out by observations.
@@ V.3.4 Small Scale Behavior
where we used $P_{a}(k)=k^{2}\,P_{\phi}(k)=1/(2\alpha_{T}k^{2})$. This is just the more precise version of the earlier estimate in Eq. (49). (Here the infrared part of the integral is soft, so we extended the integral to $k\to 0$.) We note that if we were to compute the variance in a particle’s acceleration by taking ${\bf y}\to{\bf x}$ (i.e., $L\to 0$) this diverges. Hence this theory is again not well defined. However, one can imagine a cut off $k_{\mbox{\tiny{UV}}}$ on the UV k-modes to regulate this (perhaps near the Planck scale or so).
For a collection of nearby particles, separated by a scale $L$ above the cut off, this formula tells is that their stochastic relative acceleration can be quite large. For earth based Cavendish-type tests of gravitation, the acceleration is suppressed by a factor of $M$ the mass of the source. However, in this theory there is no suppression by the mass of the source. So even though this is a contribution to the acceleration that on small distances, only rises as $a\propto 1/\sqrt{L}$, rather than Newton’s inverse square law, the fact that there is no $M$ suppression means it can be relatively large on small scales.
@@
An alternative to avoid this conclusion would be to lower the UV cut off so much that any sub-galactic fluctuations are suppressed (i.e., take $k_{\mbox{\tiny{UV}}}\sim k_{\mbox{\tiny{galaxy}}}$). But then such a theory of gravity is not useful over a wide range of scales (table-top, solar, etc) for which it is already well studied, and the UV problems of quantum gravity are not addressed at all.
@@ VI Temporal Stochasticity
The temporal dependence of the fluctuations is something that should also be carefully addressed. As mentioned earlier, the theory may conceivably lead to changes on the Hubble time or other dynamical times. In fact, as discussed earlier, the path integral suggests that the distribution should in fact jump significantly from one time step to the next, unless regulated by the matter sector. This could lead to altered constraints in the problems identified above. On the one hand, some constraints could be weakened due to temporal variation getting partially washed out through temporal averaging. On the other hand, some constraints could be be strengthened due to a new kind of, as yet unseen, temporal jitter in the behavior of gravity. All this deserves further consideration.

### v6

@@ Abstract
We analyze consequences of trying to replace dark matter and dark energy with models of stochastic spacetime. In particular, we analyze the model put forth by Ref. Oppenheim:2024rcp , in which it is claimed that “post-quantum classical gravity” (PQCG), a stochastic theory of gravity, leads to modified Newtonian dynamics (MOND) behavior on galactic scales that reproduces galactic rotation curves, and leads to dark energy. We show that this analysis has four basic problems: (i) the equations of PQCG do not lead to a new large scale force of the form claimed in the paper, (ii) the form claimed is not of the MONDian form anyhow and so does not correspond to observed galactic dynamics, (iii) the spectrum of fluctuations is very different from observations, and (iv) we also identify some theoretical problems in these models.
@@ I Introduction
An outstanding problem in modern physics is the successful unification of quantum mechanics and gravity. An interesting approach to this problem has been put forward in Refs. Oppenheim:2018igd ; Oppenheim:2023izn ; Layton:2023oud in which matter is treated quantum-mechanically, while gravity is treated classically; this is dubbed “post-quantum classical gravity” (PQCG). The coupling is such that gravity becomes effectively stochastic. Whether this framework truly leads to an internally consistent theory is not the primary focus of this paper, though we do identify some theoretical problems in this work also.
What is relevant to this paper is the possibility of such theories leading to large scale testable predictions. Very interestingly, in Ref. Oppenheim:2024rcp it was claimed that PQCG indeed does so, namely that it leads to a new long range force between matter of the modified Newtonian dynamics (MOND) form; the form that can reproduce galactic rotation curves Milgrom:1983ca ; Milgrom:1983pn ; Milgrom:1983zz ; Bekenstein:1984tv . While it would be very interesting if the unification of gravity and quantum mechanics leads to such large scale effects. Here we point out that the PQCG theory of the form presented in Oppenheim:2024rcp does not in fact lead to anything like MONDian dynamics. Furthermore, we show that the theory has a very different spectrum of fluctuations than that observed. Whether some other variation of this framework improves upon this is beyond the scope of this paper.
@@ II The Newtonian Limit of PQCG
The full PQCG theoretical framework is an interesting theory in which quantum dynamics of matter and classical dynamics of gravity are coupled together in a novel way. Nevertheless, the dynamics can be encoded in an action Oppenheim:2023izn . The full relativistic theory is somewhat complicated. However, for the purpose of studying galactic dynamics, we only need to pay attention to the low velocity limit of the theory as the characteristic speeds of gas, stars, and satellites in a galaxy are orders of magnitude slower than the speed of light (we set $c=1$). In Ref. Layton:2023oud , the low velocity limit of the theory is explained to be given by the following effective action (see Section VI for some discussion of relativistic corrections)
@@
If no boundary conditions are specified, then the most probable path simply minimizes the factor in brackets in Eq. (1). This gives the standard Newtonian potential $\Phi_{\mbox{\tiny{MPP}}}=\Phi_{N}$ obeying the Poisson equation
@@ III Solution of Modified Poisson Equation with Spherical Symmetry
For this section and the next, we work under the assumption that the fluctuations are small, leaving us with the MNE as the relevant equation for $\Phi$ assuming some non-trivial boundary conditions are imposed (then we include fluctuations in Section V). The MNE is so similar to the standard Poisson equation for Newtonian gravity that we should expect very similar behavior with only subtle differences. However, Ref. Oppenheim:2024rcp claimed that there are dramatic differences. To unpack this, let us consider a localized source and consider the region outside of the source where $\rho=0$. In this vacuum region of space, the equation reduces to
@@
With the assumption of spherical symmetry, the general solution of this equation away from $r=0$ is
@@
where $\kappa_{-1,0,1,2}$ are constants in space, although it is not obvious they should be static in time unless static boundary conditions are imposed (see below for more discussion of time dependence). By comparing to the Newtonian theory, we can easily identify
@@ III.3 Another Derivation
Then it was indicated that, apart from the constant term, the remaining $2GM\kappa_{1}/r^{2}$ term can obtain MOND. By taking a square root to recover the acceleration, this appears to give the desired $\propto\sqrt{M}/r$ force law of MONDian dynamics. And the constant pre-factor $\kappa_{1}$ is to play the role of $a_{0}/2$ of MOND, where $a_{0}\sim 10^{-10}\,\mbox{m}/\mbox{s}^{2}$ is the critical acceleration in which Newton’s law transitions from $1/r^{2}$ to $1/r$.
However, this procedure is incorrect and is the second problem in the analysis. One cannot take a sum of two terms, $1/r^{2}$ and constant, square the sum, and note that there is a cross term whose square root has a geometric mean of the desired $\sqrt{M}/r$ form.
Instead the acceleration is Newton’s $G_{N}M/r^{2}$ plus a constant. And there is no evidence that a constant correction helps to reproduce the observed galaxy rotation curves. Test particles do not respond to other objects in a way independent of their distance or mass.
@@ V.2 Full Distribution
Let us now examine in some detail the actual distribution. Often probability distributions can have interesting temporal dependence through the $dt$ integral, giving a rule for how to update the distribution from an initial time $t_{i}$ to a final time $t_{f}$, as
@@
[…]
For now, we shall proceed as is done in Ref. Oppenheim:2024rcp by ignoring the $dt$ integral and the temporal dependence. If there are no prior fixed boundary conditions, then the probability distribution for a fluctuation $\phi$ at any moment in time is the Gaussian distribution
@@
where $\alpha_{T}\sim\alpha\,\epsilon$ has units of length (we set $c=1$ here; if we reinstate factors of $c$, it has units of time4/length3).
@@ V.3 Power Spectrum
We emphasize this spectrum necessarily follows from the set-up laid out in Ref. Oppenheim:2024rcp (although the static/non-relativistic treatment is questionable and will be addressed in the next Section.) We also note that since this is all derived in kind of Newtonian approximation, we only expect it to apply on sub-horizon scales, i.e., $k\gtrsim H_{0}$ the Hubble constant.
@@
$\langle\phi({\bf x})\,\phi({\bf y})\rangle$ $=$ $\int{d^{3}k\over(2\pi)^{3}}e^{i{\bf k}\cdot({\bf x}-{\bf y})}P_{\phi}(k)$ (38)
@@
$\sigma_{k}=\sqrt{k^{3}\,P_{\phi}(k)\over 2\pi^{2}}={1\over 2\pi\sqrt{\alpha_{T}\,k}}$ (40)
@@ V.3.2 Large Scale Structure
Moreover, on sub-horizon scales, we can test if this spectrum of fluctuations is compatible with observations. The concordance model (CM) in cosmology with baryons, dark matter, and dark energy has a spectrum of fluctuations for $\Phi=\Phi_{N}$ (deterministic, but arising from $\nabla^{2}\Phi_{N}/a_{s}^{2}=4\pi G\bar{\rho}_{m}\delta_{m}=3H^{2}\Omega_{m}\delta_{m}/2$ due to inhomogeneous matter $\delta\rho_{m}=\bar{\rho}_{m}\,\delta_{m}$ drawn from some distribution). This is known to be nicely compatible with observations. In linear theory, it takes on the form (for a review, see Ref. Hertzberg:2012qn )
@@
i.e., a flat spectrum. (In the CM, $P_{m,\mbox{\tiny{CM}}}$ is physically interpreted as the “matter power spectrum”, while in PQCG it does not directly have this interpretation as the fluctuations $\phi$ have no source. But what we observe are the effects of the gravitational potential, and so this factor of ($4k^{4}/(9\Omega_{m,0}^{2}H_{0}^{4})$ can be viewed as just a convenient re-scaling. However, in this case the mean about which one is expanding only arises from the baryons, as the fluctuations $\phi$ have vanishing mean. Having the mean all come from baryons is already very strong indication that such a framework cannot match cosmological data and provide a flat universe, etc, but the details of the fluctuations are highly problematic, as we focus on here.
@@
We see that the prediction from PQCG of a flat power spectrum in the $P_{m}(k)$ variable is very different from the observed power in Figure 1. By comparing Eqs. (37) and (46), if $\alpha_{T}\sim 0.01/H_{0}$ as chosen in Ref. Oppenheim:2024rcp , then the power in PQCG is $\approx 10$ orders of magnitude too large on the scale of the horizon. And furthermore, it remains far too large even for modes that are several orders of magnitude within the horizon. This spectrum is clearly ruled out by observations.
@@ V.3.4 Small Scale Behavior
where we used $P_{a}(k)=k^{2}\,P_{\phi}(k)=1/(2\alpha_{T}k^{2})$. This is just the more precise version of the earlier estimate in Eq. (49). (Here the infrared part of the integral is soft, so we extended the integral to $k\to 0$.) We note that if we were to compute the variance in a particle’s acceleration by taking ${\bf y}\to{\bf x}$ (i.e., $L\to 0$) this diverges. Hence this theory is again not well defined. However, one can imagine a cut off $k_{\mbox{\tiny{UV}}}$ on the UV k-modes to regulate this (perhaps near the Planck scale or so). Relatedly, one can consider the physically important quantity of relative accelerations $\langle({\bf a}({\bf x})-{\bf a}({\bf y}))^{2}\rangle$ and regulate accordingly.
For a collection of nearby particles, separated by a scale $L$ above the cut off, this formula tells us that their stochastic relative acceleration can be quite large. For earth based Cavendish-type tests of gravitation, the acceleration is suppressed by a factor of $M$ the mass of the source. However, in this theory there is no suppression by the mass of the source. So even though this is a contribution to the acceleration that on small distances, only rises as $a\propto 1/\sqrt{L}$, rather than Newton’s inverse square law, the fact that there is no $M$ suppression means it can be relatively large on small scales.
@@
An alternative to avoid this conclusion would be to lower the UV cut off so much that any sub-galactic fluctuations are suppressed (i.e., take $k_{\mbox{\tiny{UV}}}\sim k_{\mbox{\tiny{galaxy}}}$). But then such a theory of gravity is not useful over a wide range of scales (table-top, solar system, etc) for which it is already well studied, and the UV problems of quantum gravity are not addressed at all.
@@ VI Temporal Stochasticity
The temporal dependence of the fluctuations is something that should also be carefully addressed. As mentioned earlier, the theory may conceivably lead to changes on the Hubble time or other dynamical times. In fact, as discussed earlier, the path integral suggests that the distribution should in fact jump significantly from one time step to the next. This could lead to altered constraints in the problems identified above. On the one hand, some constraints could be weakened due to temporal variation getting partially washed out through temporal averaging. On the other hand, some constraints could be be strengthened due to a new kind of, as yet unseen, temporal jitter in the behavior of gravity. All this deserves careful consideration.
VI.1 Relativistic Corrections
As a step in this direction, let us reinstate relativistic corrections. The full action proposed is
$\mathcal{I}=-\hat{\alpha}\int dt\,d^{3}x\sqrt{-g}\Big{[}(G_{\mu\nu}-8\pi G_{N}T_{\mu\nu})^{2}$
$\,\,\,-\beta(G-8\pi G_{N}T)^{2}\Big{]}$ (57)
where $G_{\mu\nu}$ is the Einstein tensor, $T_{\mu\nu}$ is the energy-momentum tensor, and $\hat{\alpha},\,\beta$ are constants. To study the fluctuations, we can expand around the solution of Einstein’s equations as
$g_{\mu\nu}=g_{\mu\nu,E}+h_{\mu\nu}$ (58)
where $g_{\mu\nu,E}$ obeys the Einstein field equations $G_{\mu\nu,E}=8\pi G_{N}T_{\mu\nu}$ and $h_{\mu\nu}$ is a fluctuation. Let us consider scalar fluctuations in the metric as
$h_{\mu\nu}=2\,\phi\,\delta_{\mu\nu}$ (59)
Then working to quadratic order, we obtain
[…]
where $\alpha=4(1-\beta)\hat{\alpha},\,b=3(1-3\beta)/(1-\beta)$. Here one demands $\beta<1/3$ (or the stronger constraint $\beta\leq 0$) for positivity of $b$, along with $\alpha>0$. Note that one picks up time derivatives here; such terms are in fact suppressed by factors of $1/c$ if we reinstate units by replacing $\ddot{\phi}\to\ddot{\phi}/c^{2}$. This means that when inserted into the path integral, it is better behaved in the sense that the distribution will not jump around at every instant.
By passing to Fourier space, we can compute the 2-point correlation function in time as
$\langle\phi_{k}(t)\,\phi^{*}_{k^{\prime}}(t^{\prime})\rangle=(2\pi)^{3}\delta^{3}({\bf k}-{\bf k}^{\prime})Q_{\phi}(k;t,t^{\prime})$ (61)
where the “power spectrum” $Q$, including temporal correlations, is
$Q_{\phi}(k;t,t^{\prime})=\int{d\omega\over(2\pi)}{e^{-i\omega(t-t^{\prime})}\over\alpha(k^{4}-2b\,k^{2}\,\omega^{2}/3+b\,\omega^{4})}$ (62)
We note that for $b>0$ there are no poles along the real $\omega$ line; this corresponds to correlations being exponentially suppressed in time.
This integral can be carried out using the residue theorem. The full details are not so important, but the qualitative structure is
$Q_{\phi}(k;t,t^{\prime})={e^{-a\,c\,k\,|t-t^{\prime}|}\over\alpha\,k^{3}}\,f(a\,c\,k\,|t-t^{\prime}|)$ (63)
where $a>0$ is an $\mathcal{O}(1)$ number (assuming $\mathcal{O}(1)$ values of $\beta$) and $f$ is a periodic function with period 1 and an $\mathcal{O}(1)$ amplitude. We have instated a factor of $c$ into the exponent to highlight the role that $c$ is playing here. the exponential factor shows that (up to an $\mathcal{O}(1)$ factor) there is only temporal correlation for a period of time
$T_{k}={1\over a\,c\,k}$ (64)
for the mode of interest. If we were to take the $c\to\infty$ limit, i.e., the Newtonian limit, this becomes
$Q_{\phi}(k;t,t^{\prime})\sim{1\over\alpha\,k^{4}}\delta(t-t^{\prime})$ (65)
i.e., the fluctuations become uncorrelated in time in this Newtonian limit, as we already discussed below Eq. (33). At the equal time $t=t^{\prime}$ the power spectrum
$P_{\phi}(k)=Q_{\phi}(k;t,t)$ (66)
would be formally infinite as this is where the delta-function hits. This again reinforces the points made earlier, as we defined a new parameter $\alpha_{T}\sim\alpha\,\epsilon$ where $\epsilon$ was some temporal cut off, to give the power $P\sim 1/(\alpha_{T}\,k^{4})$. These steps were needed to obtain the kind of static analysis of Ref. Oppenheim:2024rcp who ignored the temporal integral.
By keeping $\alpha$ a finite parameter of the theory and properly tracking the time dependence, we see that relativistic effects regulate the temporal correlation function then away from a delta-function to the exponential factor. We can think of this as the replacement
$\delta(t-t^{\prime})\to{a\,c\,k\over 2}e^{-a\,c\,k|t-t^{\prime}|}$ (67)
as these 2 functions have the same form in the $c\to\infty$ limit.
At equal times in the relativistic theory, we have
$P_{\phi}(k)={f(0)\over\alpha\,k^{3}}$ (68)
with $f(0)$ an $\mathcal{O}(1)$ number. So one can draw the fluctuations from this scale invariant spectrum at a given moment in time, but bearing in mind that they will deviate away from this by an $\mathcal{O}(1)$ amount on the time scale $T_{k}\sim 1/(c\,k)$, the light crossing time over a wavelength. By considering the standard deviation per log interval it is now
$\sigma_{k}=\sqrt{k^{3}\,P_{\phi}(k)\over 2\pi^{2}}\sim{0.1\over\sqrt{\alpha}}$ (69)
So to obtain a kind of dark energy like contribution, one would take
$\alpha\sim 0.01$ (70)
so that fluctuation are $\mathcal{O}(1)$ on the scale of the horizon (this is the analogue of taking $\alpha_{T}\sim 0.01/H_{0}$ in Ref. Oppenheim:2024rcp that we described earlier, when the temporal integral was ignored). For these horizon scale modes, the correlation time-scale $T_{k}\sim 1/(c\,k)$ is of order the Hubble time. So the value of this putative dark energy would fluctuate on the Hubble time, potentially changing from positive to negative, etc. Such behavior is not supported by existing data.
More problematically, this also means the spectrum is $\approx 10$ orders of magnitude too large for sub-horizon scales; compare to the observed spectrum in Eq. (46). Again this is clearly ruled out.
If one raises $\alpha$ considerably to $\alpha\sim 10^{8}$, then while a scale invariant spectrum may at first sight seem promising, it does not have the observed break in the spectrum at matter-radiation equality; compare to Figure 1 for the related variable $P_{m}=4k^{4}/(9\Omega_{m,0}^{2}H_{0}^{4})\,P_{\phi}(k)$.
On galactic scales, a new problem is that these temporal correlations mean that there is a radical change in the gravitational field on the order the light crossing time $T_{k}\sim 1/(c\,k)$. So therefore even if stars were to be orbiting in this stochastic gravitational field that Ref. Oppenheim:2024rcp proposes replaces dark matter and provides the galactic rotation curves, the stars orbits would drastically change on a light crossing time. So for example, for flattened rotation curves starting at $r=10\,$kpc, after $\approx 30,000$ years the stars would likely start orbiting in a completing different direction as the gravitational field has completely changed. This temporal change in $\Phi$ would likely disrupt halos as there would just be an incoherent mess in the gravitational potential on times longer than any light crossing times.
Finally, let us consider small scale experiments. If we consider the 2-point correlation function for acceleration of test particles, and for simplicity treat the $f$ as slow, we have
$\langle{\bf a}({\bf x},t)\cdot{\bf a}({\bf y},t^{\prime})\rangle$ $=$ $\int d\ln k\,{\sin(kL)\over kL}{k^{3}\,Q_{a}(k,t,t^{\prime})\over 2\pi^{2}}\,\,\,\,\,$ (71)
$\sim$ ${f(0)\over 2\pi^{2}\alpha(L^{2}+a^{2}\,c^{2}\,(t-t^{\prime})^{2})}$ (72)
(in fact there is an $\mathcal{O}(1)$ correction from the details of $f$, but that it is not essential here). The physically important quantity of relative accelerations $\langle({\bf a}({\bf x},t)-{\bf a}({\bf y},t^{\prime}))^{2}\rangle$ is a simple extension of the discussion here. Note that if we now consider the equal-time 2-point correlation function it rises with small $L$ as $1/L^{2}$, rather than just $1/L$ as seen earlier in Eq. (53). By again considering a table-top Cavendish experiment at $L=30$ mm, we have
$\sqrt{\langle{\bf a}\cdot{\bf a}\rangle}_{L=30\,mm}\approx{7\times 10^{17}\sqrt{f(0)}\over\sqrt{\alpha(1+\tau^{2}})}\,\mbox{m/s}^{2}$ (73)
where $\tau=a^{2}c^{2}(t-t^{\prime})^{2}/L^{2}$. At a given moment in time $\tau=0$, this suggests that to avoid being larger than the observed value of $\sim 2\times 10^{-9}$m/s2, one requires the extraordinarily large value $\alpha\gtrsim 10^{53}$. However, since the experiment takes place over a period of time much longer than the light crossing time, one should consider the factor that there is some temporal suppression through the above $\tau$ factor. However the temporal suppression here is only power law, with $\sqrt{\langle{\bf a}\cdot{\bf a}\rangle}\propto 1/\tau$ at times longer than the light crossing time. So while the residual bound on $\alpha$ may be several orders of magnitude lower than $10^{53}$ it is highly unlikely to be anywhere near the kinds of values, like $\alpha\sim 0.01$ for non-trivial cosmological consequences to be allowed.
VI.2 Other Works